OpenStudy (anonymous):

A spinner has 10 equal sections, 4 red and 6 green. Given two spins, what is the probability it lands on red at least once?

OpenStudy (anonymous):

need help please

OpenStudy (anonymous):

P(two G) = P(G) X P(G) = 0.6 X 0.6 = 0.36 So the probability of at least one red is 1 - 0.36 = 0.64 is this right

jimthompson5910 (jim_thompson5910):

that is correct, or you can say P(at least one red) = P(1 red, 1 green) + P(1 green, 1 red) + P(1 red, 1 red) P(at least one red) = P(1 red)*P(1 green) + P(1 green)*P(1 red) + P(1 red)*P(1 red) P(at least one red) = 0.4*0.6 + 0.6*0.4 + 0.4*0.4 P(at least one red) = 0.64

OpenStudy (anonymous):

thnx so much

jimthompson5910 (jim_thompson5910):

np